OpenStudy (anonymous):

Trevor, currently 22 years old, is half of his dad's age six years ago. How old is Trevor's Dad? A.) 54 B.) 40 C.) 50 D.) 44

OpenStudy (anonymous):

40

OpenStudy (izzy18970):

(22) = 1/2(d - 6) 44 = d - 6 dad's age = 50

OpenStudy (anonymous):

okay right now, trevor is half of what his dad's age was six years ago.. so take his age, multiply it by two to get 44, and then since it was 6 years ago, you must add the six years back, giving you answer c) or 50

OpenStudy (anonymous):

ok...i have another one...ok?

OpenStudy (anonymous):

17. [2.08] Which of the following is not equivalent to the formula for the area of a rectangle A = LW? A.) A/L=W B.) A/W=L C.) A=L/W D.) A/WL=1

OpenStudy (tkhunny):

Open letter to @Izzy18970 and @LeonSquishy Don't do that. Please read the Code of Conduct.

OpenStudy (izzy18970):

don't do what??

OpenStudy (izzy18970):

ok